please help me on this word problem:
The sum of a number and 17 more than twice the same number is 101. Find the number.

Answer by CharStar(110)

1st number = x
2nd number = 17 + 2x
Total = 101
equation is:
x + 17 + 2x = 101
3x + 17 = 101
subtract 17 from both sides
3x = 84
x = 28
1st number = 28
2nd number = 17 + 2(28) 17 + 56 = 73
Total = 28 + 73 = 101